OpenStudy (anonymous):

3 over x equals x over 32 help

OpenStudy (ivettef365):

that sounds like a proportional problem

OpenStudy (ivettef365):

cross multiply

OpenStudy (zzr0ck3r):

\[\frac{3}{x}=\frac{x}{32}\] \[x^2 = 96\]

OpenStudy (anonymous):

I got 96 when I cross multiplied. that's one of my answer choices but there are also radical answers

OpenStudy (zzr0ck3r):

\[x=\pm \sqrt{96}\]

OpenStudy (anonymous):

I got it wrong :/ im gonna post the problem so you can see it

OpenStudy (anonymous):

Solve for x. 3 over x equals x over 32 96 48 4 radical 6 6 radical 4

OpenStudy (zzr0ck3r):

\[x=\pm \sqrt{48*2}=\pm \sqrt{4*4*6}=\pm 4\sqrt{6}\]

OpenStudy (anonymous):

how did you do it?

OpenStudy (zzr0ck3r):

hmm I showed the steps. Which part?

OpenStudy (anonymous):

oh wait I get it now.. Im just confused on how you got 48?

OpenStudy (zzr0ck3r):

(3/x) = (x/32) multiply each side by 32x x^2 = 96 sqrt both sides x=sqrt(96) 96 = 4*4*6 so x = sqrt(4*4*6) x = 4sqrt(6)

OpenStudy (anonymous):

ohhhh, I get it now, its not that hard. sorry to bother you. THANK YOU A LOT :)